# A Survey of Lottery Ticket Hypothesis

### 07 Mar 2024 v1

@@ 1 Introduction
The Lottery Ticket Hypothesis (LTH) (Frankle & Carbin, 2018) states that a dense neural network model contains a highly sparse subnetwork (i.e., winning tickets) that can achieve even better performance than the original model. The winning tickets can be identified by training a network and pruning its parameters with the smallest magnitude in an iterative way or one-shot way. Before it is trained in each iteration, the weights will be reset to the original initialization. Instead of resetting to the original initialization, some works (Mikler, 2021; Renda et al., 2020a; Gadhikar & Burkholz, 2024) rewind the parameters of the subnetwork to the early stage of training. The identified winning tickets are retrainable to reduce the high memory cost and long inference time of the original neural networks and the existence of winning tickets has been verified in both experiments and theory (Frankle et al., 2019; Diffenderfer & Kailkhura, 2021; Savarese et al., 2020; Zhou et al., 2019; Malach et al., 2020; Zhang et al., 2021a; Ma et al., 2021a; Zhou et al., 2019; Chen et al., 2021d; Su et al., 2020; Chen et al., 2022a). Later works have extended LTH to find the winning tickets for different kinds of neural networks such as generative models (Chen et al., 2021h; Kalibhat et al., 2021; Jiang et al., 2023), Transformers (Brix et al., 2020; Prasanna et al., 2020; Chen et al., 2020a; Behnke & Heafield, 2020), and GNNs (Chen et al., 2021c; Hui et al., 2023; Zhang et al., 2024; Harn et al., 2022) Despite the success of LTH, the iterative train-prune-retrain process and even the subsequent fine-tuning are costly to compute (Liu et al., 2021). Many works have proposed algorithms to find lottery tickets efficiently  (Yin et al., 2023a; Zhang et al., 2021c; Wang et al., 2020b; Tanaka et al., 2020; Ye et al., 2020; Evci et al., 2020; Hou et al., 2022; Ma et al., 2022; Jaiswal et al., 2023a). LTH is also related to many other domains including robustness, transfer learning, fairness, and federated learning (Yuan et al., 2021; Goli & Aamodt, 2020; Raihan & Aamodt, 2020; Chen et al., 2021e; Mehta, 2019; Morcos et al., 2019b; Itahara et al., 2020; Seo et al., 2021; Mugunthan et al., 2022; Shi et al., 2023; Hansen & Søgaard, 2021). This survey aims to provide a comprehensive overview of the state of LTH research, categorizing various approaches, highlighting open issues, and presenting a benchmark.
@@ 2 Theory
The existence of a winning ticket has been proved theoretically in many works. The first theoretical evidence of the existence of winning tickets was given by Malach et al. (2020). By approximating ReLU networks, This work proves that optimizing the weights is equivalent to pruning entire neurons of a sufficiently large random network. For every bounded distribution and every target network with bounded weights, a sufficiently over-parameterized neural network with random weights contains a subnetwork with roughly the same accuracy as the target network, without any further training. However, The polynomial over-parameterization requirement in the proof is at odds with recent experimental results with networks that are wider than the bound. An exponential improvement to the over-parameterization requirement is offered in  Pensia et al. (2020). Instead of providing significantly tighter bounds,  (Orseau et al., 2020) points out that the overparameterized network only needs a logarithmic factor number of neurons per weight of the target subnetwork. Any network with a depth of $L$ and width of $W$ can be obtained by pruning a randomly initialized network with a depth of $2L$ and width increased by a logarithmic factor, or a network with a depth of $L+1$ and width increased by another logarithmic factorBurkholz (2022b). This discovery implies that effective lottery tickets can be identified at frequently used depths, necessitating only logarithmic levels of overparameterization.
@@
Elastic LTH Chen et al. (2021f) NeurIPS Pytorch https://github.com/VITA-Group/ElasticLTH
@@
Sanity ChecksMa et al. (2021a) NeurIPS Pytorch https://github.com/boone891214/sanity-check-LTH
@@
LTH-BERTPrasanna et al. (2020) EMNLP Pytorch https://github.com/sai-prasanna/bert-experiments BERT GLUE
SubsetSumPensia et al. (2020) NeurIPS Pytorch https://github.com/acnagle/optimal-lottery-tickets LeNet5 MNIST
@@ 3.1 Graph Neural networks
Previous GLT methods use the same topology structure across all layers. The strategy hinders the integration of GLT into deeper and larger-scale GNN contexts. To address this issue, an Adaptive, Dynamic, and Automated framework Zhang et al. (2024) is introduced for identifying Graph Lottery Tickets (AdaGLT). The proposed method tailors layer-adaptive sparse structures for various datasets and GNNs and captures GLT across diverse sparsity levels. The theoretical proofs to mitigate over-smoothing issues and obtain improved sparse structures in deep GNNs are also provided.
@@ 3.2 Transformer
Large pre-trained transformers have achieved great success in the past few years. However, Due to the exponentially increasing parameter counts of these large pre-trained models, fine-tuning these models with non-industry standard hardware is becoming seemingly impossible. (Yu et al., 2019) conduct a comprehensive investigation into the applicability of the Lottery Ticket Hypothesis (LTH) within the domains of Natural Language Processing (NLP). They scrutinize recurrent LSTM models as well as large-scale Transformer models. Notably, the authors successfully identify initializations that qualified as winning tickets, facilitating Transformer models to achieve performance levels nearly equivalent to their full-size counterparts, despite a reduction in size by one-third. Two studies have addressed the feasibility of applying the LTH to BERT. (Prasanna et al., 2020) successfully identified subnetworks within fine-tuned BERT that exhibited performance on par with the complete model. They observe that similarly-sized subnetworks sampled from other parts of the model demonstrated worse performance. However, they noted that even the least performing subnetworks exhibited notable trainability, indicating the substantial utility of the majority of pre-trained BERT weights. In a separate investigation, (Chen et al., 2020a) examined pre-trained BERT and unearthed the presence of trainable and transferable subnetworks within it. Notably, these subnetworks are discovered during the pre-training phase, distinguishing the earlier research, which obtained subnetworks during the fine-tuning phase, a process necessitating a certain degree of additional training. In addition to its validation in the field of NLP, (Gan et al., 2022) explore whether trainable subnetworks, as suggested by the LTH, exist in vision-and-language models. The findings indicate that while exact matches are challenging, "relaxed" winning tickets at 50%-70% sparsity maintain 99% accuracy. Task-specific pruning leads to reasonably transferable subnetworks, and pre-training tasks at 60%/70% sparsity achieve 98%/96% accuracy. (Behnke & Heafield, 2020) investigate the pruning of attention heads in Transformers, prompted by recent research indicating a lack of confidence in the decisions made by a significant portion of attention heads. Through experiments in machine translation, they demonstrate the potential to remove up to three-quarters of attention heads from transformer-big during early training.
@@
During the fine-tuning stage of Transformer models applied to NLP tasks, the integration of adapter modules has been found to expedite transfer learning. However, adapter-fusion models, which incorporate knowledge from various tasks, can lead to redundancy and increased computational costs. To mitigate this, (Wu et al., 2022) propose a method to assess the impact of each adapter module and employ the Lottery Ticket Hypothesis to prune them effectively. (Gong et al., 2022) explore the applicability of the Lottery Ticket Hypothesis to pre-trained language models (PLMs) for fine-tuning. The authors introduce a regularization approach using L1 distance and identify a subnetwork structure termed the "dominant winning ticket." They empirically demonstrate that this dominant winning ticket can achieve performance comparable to the full-parameter model, is transferable across different tasks, and exhibits a natural structure within each parameter matrix. To address the computational cost of repetitive training and pruning routine of the transformer, Instant Soup Pruning (ISP) (Jaiswal et al., 2023b)is introduced as an efficient alternative to the iterative magnitude pruning (IMP) method. ISP leverages the concept of model soups, merging fine-tuned weights from multiple models to achieve better minima. The proposed approach replaces the costly intermediate stages of IMP with a computationally efficient weak mask generation and aggregation process. Specifically, ISP generates multiple weak and noisy subnetworks through a few iterations with varying training protocols and data subsets, then combines them to create a high-quality denoised subnetwork.
@@ 3.3 Generative Models
While prior research mainly focused on supervised learning scenarios, (Kalibhat et al., 2021) extends the concept to deep generative models including GANs and VAEs. The study demonstrates that iterative magnitude pruning, combined with generative losses, effectively identifies winning tickets, achieving high sparsity levels. Moreover, the transferability of winning tickets across different generative models implies they have inherent biases beneficial for training various deep generative models. Furthermore, the paper introduces the concept of "early-bird tickets," enabling substantial reductions in floating-point operations and training time, enhancing the feasibility of training large-scale generative models under resource constraints.
@@
An extensive exploration of the applicability of pruning techniques in generative models is conducted to Yeo et al. (2023) discover subnetworks that perform similarly to or better than the trained dense model. They introduce the concepts of weak lottery tickets and strong lottery tickets. The former refers to tickets identified from networks with learned weights, while the latter denotes tickets identified without any weight updates. The experimental result shows that a subnetwork of GFMN with only 10% of the weights remaining can reach similar FID scores as GFMN. LTH has also been applied to diffusion models (Jiang et al., 2023). It shows that we can find a winning ticket of DDPM Ho et al. (2020) at sparsity $90\%-99\%$ without compromising performance for denoising diffusion probabilistic models on benchmarks (Cifar-10, Cifar-100, MNIST). Different from existing LTH works which identify the subnetworks with a unified sparsity along different layers, this work proposes to find the winning ticket with varying sparsity along different layers in the model.
@@ 4.3 Key factors
There are three critical components in LTH: Zeros, Signs, and the Supermask Zhou et al. (2019). Three questions are answered by ablating these factors: why setting weights to zero is important, how signs are all you need to make the reinitialized network train, and why masking behaves like training. It is the sign in the original initialization, not the magnitude of the weights, that is crucial to the performance of LT networks. Also, the masking procedure can be considered as a training operation. Supermasks are observed to produce partially working networks without training. (Evci et al., 2022) emphasize that training unstructured sparse Neural Networks (NNs) from random initialization often results in poor generalization, with Lottery Tickets and Dynamic Sparse Training being notable exceptions. They observe that in the initialization phase of sparse NNs, gradient flow tends to be weak. In comparison to other methods, Dynamic Sparse Training significantly bolsters gradient flow during the training phase, potentially contributing to its effectiveness. Conversely, Lottery Tickets does not exhibit a significant improvement in gradient flow. Therefore, the authors posit three key factors contributing to the success of Lottery Tickets: 1. Proximity to the pruning solution during initialization. 2. Residing within the basin of the pruning solution. 3. Learning functions akin to the pruning solution.
@@ 5.2 Unstructured pruning
Most of the existing works in the literature focus on finding unstructured winning tickets where the positions of zeros are not predefined. These methods pruned individual elements with the smallest magnitude. (Shang et al., 2022) delve into the Magnitude-Based Pruning (MBP) scheme and provides a fresh perspective by employing Fourier analysis on the deep learning model to guide model selection. In addition to elucidating the generalization capability of MBP using the Fourier transform, a novel two-stage pruning approach is proposed. The first stage involves obtaining the topological structure of the pruned network, followed by a second stage of retraining to restore capacity using knowledge distillation from lower to higher frequencies.
Different from these methods based on IMP, (You et al., 2022a) present an approach addressing Deep Neural Networks (DNNs) termed "Two-in-One Training," which enables the direct identification of efficient DNNs and their associated lottery subnetworks, referred to as SuperTickets, from a supernet. Previous methods typically employ Neural Architecture Search (NAS) followed by pruning algorithms to obtain lottery subnetworks for efficient DNNs. This approach, known as "First-search-then-prune (S+P)," involves an initial search for an efficient network architecture and subsequent sparsity induction. In contrast, this paper proposes a unified approach that simultaneously conducts architecture search and pruning during the training of the supernet within NAS. In the search phase, the authors compute importance factors for search units and set a threshold $\epsilon$ to remove unnecessary units. In the pruning phase, they employ the Importance Masking Pruning (IMP) method to achieve the desired level of sparsity. Additionally, two techniques are introduced in the search and pruning phases. In the search phase, Iterative Reactivation is implemented. Unlike the conventional IMP method, the authors incorporate reactivation during the search process rather than after pruning. This choice is made to retain training opportunities for the sparsified network, which is believed to contribute to performance enhancement. Experimental results validate this hypothesis by comparing the performance of networks with reactivation during search versus pruning (Schotthöfer et al., 2022) introduce a novel approach to identify efficient low-rank subnetworks. These subnetworks are determined and adjusted during the training process, significantly reducing the time and memory resources required for both training and evaluation. The key concept involves restricting weight matrices to a low-rank manifold and updating the low-rank factors instead of the entire matrix during training. This method ensures that the training updates stay within the specified manifold, providing guarantees of approximation, stability, and descent.
@@
Due to the discontinuity of the step function $H$ at certain points, and its derivative being zero at all points, dealing with $H$ in computer programs poses some challenges. The authors propose an approximation method, using an S-shaped function parametrized by $\beta$ to approximate $H$. The form of this approximating function is $s\mapsto\sigma(\beta s)$, where $\sigma$ is the sigmoid function. Thus, the final objective function is:

### 12 Mar 2024 v2

@@ 1 Introduction
The Lottery Ticket Hypothesis (LTH) (Frankle & Carbin, 2018) states that a dense neural network model contains a highly sparse subnetwork (i.e., winning tickets) that can achieve even better performance than the original model. The winning tickets can be identified by training a network and pruning its parameters with the smallest magnitude in an iterative way or one-shot way. Before it is trained in each iteration, the weights will be reset to the original initialization. Instead of resetting to the original initialization, some works (Mikler, 2021; Renda et al., 2020a; Gadhikar & Burkholz, 2024) rewind the parameters of the subnetwork to the early stage of training. The identified winning tickets are retrainable to reduce the high memory cost and long inference time of the original neural networks, and the existence of winning tickets has been verified in both experiments and theory (Frankle et al., 2019; Diffenderfer & Kailkhura, 2021; Savarese et al., 2020; Zhou et al., 2019; Malach et al., 2020; Zhang et al., 2021a; Ma et al., 2021a; Zhou et al., 2019; Chen et al., 2021d; Su et al., 2020; Chen et al., 2022a). Later works have extended LTH to find the winning tickets for different kinds of neural networks such as generative models (Chen et al., 2021h; Kalibhat et al., 2021; Jiang et al., 2023), Transformers (Brix et al., 2020; Prasanna et al., 2020; Chen et al., 2020a; Behnke & Heafield, 2020), and GNNs (Chen et al., 2021c; Hui et al., 2023; Zhang et al., 2024; Harn et al., 2022) Despite the success of LTH, the iterative train-prune-retrain process and even the subsequent fine-tuning are costly to compute (Liu et al., 2021). Many works have proposed algorithms to find lottery tickets efficiently  (Yin et al., 2023a; Zhang et al., 2021c; Wang et al., 2020b; Tanaka et al., 2020; Ye et al., 2020; Evci et al., 2020; Hou et al., 2022; Ma et al., 2022; Jaiswal et al., 2023a). LTH is also related to many other domains including robustness, transfer learning, fairness, and federated learning (Yuan et al., 2021; Goli & Aamodt, 2020; Raihan & Aamodt, 2020; Chen et al., 2021e; Mehta, 2019; Morcos et al., 2019b; Itahara et al., 2020; Seo et al., 2021; Mugunthan et al., 2022; Shi et al., 2023; Hansen & Søgaard, 2021). This survey aims to provide a comprehensive overview of the state of LTH research, categorizing various approaches, highlighting open issues, and presenting a benchmark.
@@ 2 Theory
The existence of a winning ticket has been proved theoretically in many works. The first theoretical evidence of the existence of winning tickets was given by Malach et al. (2020). By approximating ReLU networks, This work proves that optimizing the weights is equivalent to pruning entire neurons of a sufficiently large random network. For every bounded distribution and every target network with bounded weights, a sufficiently over-parameterized neural network with random weights contains a subnetwork with roughly the same accuracy as the target network, without any further training. However, the polynomial over-parameterization requirement in the proof is at odds with recent experimental results with networks that are wider than the bound. An exponential improvement to the over-parameterization requirement is offered in  Pensia et al. (2020). Instead of providing significantly tighter bounds,  (Orseau et al., 2020) points out that the overparameterized network only needs a logarithmic factor number of neurons per weight of the target subnetwork. Any network with a depth of $L$ and width of $W$ can be obtained by pruning a randomly initialized network with a depth of $2L$ and width increased by a logarithmic factor, or a network with a depth of $L+1$ and width increased by another logarithmic factor (Burkholz, 2022b). This discovery implies that effective lottery tickets can be identified at frequently used depths, necessitating only logarithmic levels of overparameterization.
@@
Elastic LTH (Chen et al., 2021f) NeurIPS Pytorch https://github.com/VITA-Group/ElasticLTH
@@
Sanity Checks (Ma et al., 2021a) NeurIPS Pytorch https://github.com/boone891214/sanity-check-LTH
@@
LTH-BERT (Prasanna et al., 2020) EMNLP Pytorch https://github.com/sai-prasanna/bert-experiments BERT GLUE
SubsetSum (Pensia et al., 2020) NeurIPS Pytorch https://github.com/acnagle/optimal-lottery-tickets LeNet5 MNIST
@@ 3.1 Graph Neural networks
Previous GLT methods use the same topology structure across all layers. The strategy hinders the integration of GLT into deeper and larger-scale GNN contexts. To address this issue, an Adaptive, Dynamic, and Automated framework (Zhang et al., 2024) is introduced for identifying Graph Lottery Tickets (AdaGLT). The proposed method tailors layer-adaptive sparse structures for various datasets and GNNs and captures GLT across diverse sparsity levels. The theoretical proofs to mitigate over-smoothing issues and obtain improved sparse structures in deep GNNs are also provided.
@@ 3.2 Transformer
Large pre-trained transformers have achieved great success in the past few years. However, due to the exponentially increasing parameter counts of these large pre-trained models, fine-tuning these models with non-industry standard hardware is becoming seemingly impossible. (Yu et al., 2019) conduct a comprehensive investigation into the applicability of the Lottery Ticket Hypothesis (LTH) within the domains of Natural Language Processing (NLP). They scrutinize recurrent LSTM models as well as large-scale Transformer models. Notably, the authors successfully identify initializations that qualified as winning tickets, facilitating Transformer models to achieve performance levels nearly equivalent to their full-size counterparts, despite a reduction in size by one-third. Two studies have addressed the feasibility of applying the LTH to BERT. (Prasanna et al., 2020) successfully identified subnetworks within fine-tuned BERT that exhibited performance on par with the complete model. They observed that similarly sized subnetworks sampled from other parts of the model demonstrated worse performance. However, they noted that even the least performing subnetworks exhibited notable trainability, indicating the substantial utility of the majority of pre-trained BERT weights. In a separate investigation, (Chen et al., 2020a) examined pre-trained BERT and unearthed the presence of trainable and transferable subnetworks within it. Notably, these subnetworks were discovered during the pre-training phase, distinguishing the earlier research, which obtained subnetworks during the fine-tuning phase, a process necessitating a certain degree of additional training. In addition to its validation in the field of NLP, (Gan et al., 2022) explore whether trainable subnetworks, as suggested by the LTH, exist in vision-and-language models. The findings indicate that while exact matches are challenging, "relaxed" winning tickets at 50%-70% sparsity maintain 99% accuracy. Task-specific pruning leads to reasonably transferable subnetworks, and pre-training tasks at 60%/70% sparsity achieve 98%/96% accuracy. (Behnke & Heafield, 2020) investigate the pruning of attention heads in Transformers, prompted by recent research indicating a lack of confidence in the decisions made by a significant portion of attention heads. Through experiments in machine translation, they demonstrate the potential to remove up to three-quarters of attention heads from transformer-big during early training.
@@
During the fine-tuning stage of Transformer models applied to NLP tasks, the integration of adapter modules has been found to expedite transfer learning. However, adapter fusion models, which incorporate knowledge from various tasks, can lead to redundancy and increased computational costs. To mitigate this, (Wu et al., 2022) propose a method to assess the impact of each adapter module and employ the Lottery Ticket Hypothesis to prune them effectively. (Gong et al., 2022) explore the applicability of the Lottery Ticket Hypothesis to pre-trained language models (PLMs) for fine-tuning. The authors introduce a regularization approach using L1 distance and identify a subnetwork structure termed the "dominant winning ticket." They empirically demonstrate that this dominant winning ticket can achieve performance comparable to the full-parameter model, is transferable across different tasks, and exhibits a natural structure within each parameter matrix. To address the computational cost of repetitive training and pruning routine of the transformer, Instant Soup Pruning (ISP) (Jaiswal et al., 2023b) is introduced as an efficient alternative to the iterative magnitude pruning (IMP) method. ISP leverages the concept of model soups, merging fine-tuned weights from multiple models to achieve better minima. The proposed approach replaces the costly intermediate stages of IMP with a computationally efficient weak mask generation and aggregation process. Specifically, ISP generates multiple weak and noisy subnetworks through a few iterations with varying training protocols and data subsets, then combines them to create a high-quality denoised subnetwork.
@@ 3.3 Generative Models
While prior research mainly focused on supervised learning scenarios, (Kalibhat et al., 2021) extend the concept to deep generative models including GANs and VAEs. The study demonstrates that iterative magnitude pruning, combined with generative losses, effectively identifies winning tickets, achieving high sparsity levels. Moreover, the transferability of winning tickets across different generative models implies they have inherent biases beneficial for training various deep generative models. Furthermore, the paper introduces the concept of "early-bird tickets," enabling substantial reductions in floating-point operations and training time, enhancing the feasibility of training large-scale generative models under resource constraints.
@@
An extensive exploration of the applicability of pruning techniques in generative models is conducted by Yeo et al. (2023) to discover subnetworks that perform similarly to or better than the trained dense model. They introduce the concepts of weak lottery tickets and strong lottery tickets. The former refers to tickets identified from networks with learned weights, while the latter denotes tickets identified without any weight updates. The experimental result shows that a subnetwork of GFMN with only 10% of the weights remaining can reach similar FID scores as GFMN. LTH has also been applied to diffusion models (Jiang et al., 2023). It shows that we can find a winning ticket of DDPM (Ho et al., 2020) at sparsity $90\%-99\%$ without compromising performance for denoising diffusion probabilistic models on benchmarks (Cifar-10, Cifar-100, MNIST). Different from existing LTH works which identify the subnetworks with a unified sparsity along different layers, this work proposes to find the winning ticket with varying sparsity along different layers in the model.
@@ 4.3 Key factors
There are three critical components in LTH: Zeros, Signs, and the Supermask (Zhou et al., 2019). Three questions are answered by ablating these factors: why setting weights to zero is important, how signs are all you need to make the reinitialized network train, and why masking behaves like training. It is the sign in the original initialization, not the magnitude of the weights, that is crucial to the performance of LT networks. Also, the masking procedure can be considered as a training operation. Supermasks are observed to produce partially working networks without training. (Evci et al., 2022) emphasize that training unstructured sparse Neural Networks (NNs) from random initialization often results in poor generalization, with Lottery Tickets and Dynamic Sparse Training being notable exceptions. They observe that in the initialization phase of sparse NNs, gradient flow tends to be weak. In comparison to other methods, Dynamic Sparse Training significantly bolsters gradient flow during the training phase, potentially contributing to its effectiveness. Conversely, Lottery Tickets does not exhibit a significant improvement in gradient flow. Therefore, the authors posit three key factors contributing to the success of Lottery Tickets: 1. Proximity to the pruning solution during initialization. 2. Residing within the basin of the pruning solution. 3. Learning functions akin to the pruning solution.
@@ 5.2 Unstructured pruning
Most of the existing works in the literature focus on finding unstructured winning tickets where the positions of zeros are not predefined. These methods pruned individual elements with the smallest magnitude. (Shang et al., 2022) delve into the Magnitude-Based Pruning (MBP) scheme and provide a fresh perspective by employing Fourier analysis on the deep learning model to guide model selection. In addition to elucidating the generalization capability of MBP using the Fourier transform, a novel two-stage pruning approach is proposed. The first stage involves obtaining the topological structure of the pruned network, followed by a second stage of retraining to restore capacity using knowledge distillation from lower to higher frequencies.
Different from these methods based on IMP, (You et al., 2022a) present an approach addressing Deep Neural Networks (DNNs) termed "Two-in-One Training," which enables the direct identification of efficient DNNs and their associated lottery subnetworks, referred to as SuperTickets, from a supernet. Previous methods typically employ Neural Architecture Search (NAS) followed by pruning algorithms to obtain lottery subnetworks for efficient DNNs. This approach, known as "First-search-then-prune (S+P)," involves an initial search for an efficient network architecture and subsequent sparsity induction. In contrast, this paper proposes a unified approach that simultaneously conducts architecture search and pruning during the training of the supernet within NAS. In the search phase, the authors compute importance factors for search units and set a threshold $\epsilon$ to remove unnecessary units. In the pruning phase, they employ the Importance Masking Pruning (IMP) method to achieve the desired level of sparsity. Additionally, two techniques are introduced in the search and pruning phases. In the search phase, Iterative Reactivation is implemented. Unlike the conventional IMP method, the authors incorporate reactivation during the search process rather than after pruning. This choice is made to retain training opportunities for the sparsified network, which is believed to contribute to performance enhancement. Experimental results validate this hypothesis by comparing the performance of networks with reactivation during search versus pruning. Schotthöfer et al. (2022) introduce a novel approach to identify efficient low-rank subnetworks. These subnetworks are determined and adjusted during the training process, significantly reducing the time and memory resources required for both training and evaluation. The key concept involves restricting weight matrices to a low-rank manifold and updating the low-rank factors instead of the entire matrix during training. This method ensures that the training updates stay within the specified manifold, providing guarantees of approximation, stability, and descent.
@@
Due to the discontinuity of the step function $H$ at certain points, and its derivative being zero at all points, dealing with $H$ in computer programs poses some challenges. The authors propose an approximation method, using an S-shaped function parameterized by $\beta$ to approximate $H$. The form of this approximating function is $s\mapsto\sigma(\beta s)$, where $\sigma$ is the sigmoid function. Thus, the final objective function is: